x(y^2+z)p-y(x^2+z)q=z(x^2-y^2)


1
Expert's answer
2020-08-11T18:21:14-0400
"Solution"

Given


"x(y^2+z)p-y(x^2+z)q=z(x^2-y^2)--------->(1)"

Lagrange's auxiliary equation of (1) are;


"\frac{dx}{x(y^2+z)}=\frac{dy}{-y(x^2+z)}=\frac{dz}{z(x^2-y^2)}"

Choosing "(\frac{1}{x}, \frac{1}{y}, \frac{1}{z})" as multipliers, we get


"\frac{dx}{x}, \frac{dy}{y}, \frac{dz}{z}=0"

Integrating "\frac{dx}{x}+ \frac{dy}{y}+\frac{dz}{z}=0" we get "xyz=c_1"


Choosing "(x,y,-1)" as multipliers, we have


"\frac{xdx+ydy-dz}{x^2y^2+x^2z-y^2x^2-y^2z-x^2z+y^2z}=\frac{xdx+ydy-dz}{0}"


Integrating "xdx+ydy-dz=0"


"\frac{x^2}{2}+ \frac{y^2}{2}-z=c \implies x^2+y^2-2z=c_2"

"\therefore" the solution is the system of equations:


"xyz=c_1"

"x^2+y^2-2z=c_2"

Taking "t" as a parameter, the given equation of the straight line "x+y=0,z=1" can be put in parametric form "x=t,y=-t,z=1"


Using this,


"xyz=c_1"

"x^2+y^2-2z=c_2"

May be written as,


"-t^2=c_1"

"2t^2-2=c_2"

Eliminating "t" from the equation we get,


"2(-c_1)-2=c_2 \implies 2c_1+c_2+2=0"

Putting values of "c_1" and "c_2" , the desired integral surface is,


"2xyz+x^2+y^2-2z+2=0------>Answer"